What is the pH of a buffer that was prepared by adding 3.96 g of sodium benzoate, NaC7H5O2, to 1.00 L of 0.0100 M benzoic acid, HC7H5O2? Assume that there is no change in volume. The K a for benzoic acid is 6.3 × 10–5.
Answer #1

pKa = -log Ka = -log(6.3*10^-5) = 4.2

moles of sodium benzoate = 3.96 gm/144.11g/mol = 0.0275 mol

moles of benzoic acid = 1L * 0.01M = 0.01 moles

pH = pKa + Log[benzoate/benzoic acid]

     = 4.2 + log [0.0275/0.01] = 4.64

ph of the solution is 4.64